Calculus I, Fall 2015 (http://www.math.nthu.edu.tw/˜ wangwc/)

Brief answer to selected problems in HW14

1. Chap 6, Additional and Advanced Exercises: problems 4 Left hand side = arclength of the curve y = sin x from (0, 0) to (α, sin α). Right hand side is the distance between (0, 0) to (α, sin α).